# ConDiSim: Conditional Diffusion Models for Simulation-Based Inference

### v1

@@ 4 Experiments
We evaluate ConDiSim on 10 benchmark problems from (sbibm) suite [Lueckmann et al., 2021] and two real-world problems. For each benchmark, we train with simulation budgets of 10 0001000010\,00010 000, 20 0002000020\,00020 000, and 30 0003000030\,00030 000 simulations, and infer the posterior by drawing 10 0001000010\,00010 000 samples from the model. Posterior quality is assessed with Maximum Mean Discrepancy (MMD), the Classifier Two-Sample Test (C2ST), and Empirical Cumulative Distribution Function (ECDF) diagnostics implemented in BayesFlow [Radev et al., 2023b]. ECDF comparisons were performed using the BayesFlow framework [Radev et al., 2023b]; further details and ECDF plots for various benchmarks are provided in Appendix E. Complete hyperparameter configurations for all methods are detailed in Appendix B.
@@
(a) SLCP
(b) SLCP Distractors
Figure 3: SLCP: Comparison of generated vs. reference posteriors for SLCP and SLCP Distractors
4.1 Benchmark Posterior Analysis
In this section, we analyze three representative benchmarks—Two Moons, SLCP, and SLCP Distractors—selected for their distinct challenges in posterior inference. The Two Moons benchmark poses the challenge of capturing a bi-modal posterior with a distinctive crescent geometry. As depicted in Figure 2(a), ConDiSim accurately recovers both modes and the complex posterior structure. The ECDF plots for $\theta_{1}$ and $\theta_{2}$ (Figure 2(b)) demonstrate that the inferred distributions are well-calibrated, with the majority of rank statistics falling within the 90% confidence bands; a minor deviation for $\theta_{1}$ suggests a slight degree of overconfidence.
The SLCP benchmark encompasses a range of parameter behaviors - including uniform, Gaussian, multi-modal, and skewed distributions. Its variant, SLCP Distractors, further tests the model’s robustness by incorporating uninformative noise dimensions, thereby simulating scenarios with heavily corrupted data. Figures 3(a) and 3(b) illustrate that ConDiSim is able to capture these complex posterior landscapes with high fidelity, and also effectively isolate the informative parameters from the distractors. The rank ECDF plots for both SLCP variants (Figures 14(a) and 14(b) in the Appendix) also remain centered within the confidence intervals, confirming that the model reliably captures parameter variability even under noisy conditions. Detailed descriptions of these benchmarks, are provided in Appendix A.
Table 1 presents the C2ST scores for the 10 benchmarking problems. For the Two Moons task, ConDiSim achieves C2ST performance comparable to Simformer, APT, and NPE, while requiring the shortest training time among the evaluated models (see Table 2). Simformer and APT perform well on most test problems – Simformer leveraging its transformer architecture with cross-attention, while APT is able to iteratively refine the posterior approximation. The SLCP problem remains particularly challenging, with all methods struggling to accurately infer the true posterior. It is important to note that the C2ST metric is sensitive to parameter estimation errors and can be difficult to interpret in high-dimensional settings, as a single poorly estimated dimension may disproportionately affect the overall score. For instance, as illustrated in Figure 3(a), ConDiSim encounters difficulties with the final dimension, despite accurately capturing the other dimensions, which is reflected in an overall C2ST score of 0.82. To provide additional insight, we report the MMD metric for each benchmark problem in the Supplementary Table 4.
@@ 4.1 Benchmark Posterior Analysis
Figure 4: Posterior Distributions for Hodgkin–Huxley model
@@
(b) Posterior Predictive Trajectories
Figure 5: Posterior Distributions for Genetic Oscillator model
@@ 4.2 Application to Real-World Scenarios
We evaluate the performance of ConDiSim on two real-world simulation models. First, the Hodgkin-Huxley (HH) model [Gloeckler et al., 2024] which simulates neuronal action potentials by modeling the membrane voltage dynamics governed by sodium, potassium, and leak currents. Second, we consider the Genetic Oscillator model which is a high-dimensional gene regulatory network that exhibits oscillatory behavior, driven by a positive-negative feedback loop mimicking a circadian clock [Vilar et al., 2002]. The model encompasses 9 species undergoing 18 reactions, parameterized by 15 reaction rate constants. Although the full system involves multiple interacting components, we focus on the inference of only three species (C, A, R). Detailed description is available in Appendix A.
Unlike the SBI benchmarking problems, the true posterior distributions for these real-world models are unknown. To assess ConDiSim, we use well motivated parameter combinations from literature to generate corresponding synthetic observations (e.g., parameter combination resulting in stable oscillations in case of the genetic oscillator). These observations serve as the observed data for inference, enabling us to evaluate the inferred posterior by comparing its concentration around the true parameter value. We also perform a posterior predictive check to validate the model. In this procedure, five posterior samples are randomly selected and used to simulate corresponding observations, which are then compared with the observed data. This check provides an additional measure of the consistency and reliability of the inferred posteriors.
For the HH model, Figure 4(a) shows that ConDiSim effectively infers a posterior distribution concentrated around the true parameter values, while Figure 4(b) confirms that the simulated voltage trajectories closely match the reference signals. The parameter estimation task for the genetic oscillator is more challenging due to broad and varied parameter priors. Figure 5(a) indicates that the peak of the posterior distribution is near the true parameter values. Figure 5(b) illustrates both the true and generated trajectories for three selected species, obtained by randomly sampling three posterior sets and plotting the corresponding simulations. Initially, these trajectories align well with the true data, with slight deviations emerging over time due to the inherent stochastic nature of the model. Figure 7 (Appendix) depicts multiple trajectories simulated using the true parameter combination - which closely match the inferred results. In both cases, the trajectories are very close to each other during initial time steps, and develop small deviations owing to model stochasticity as time steps evolve. The plots highlight the system’s inherent variability, and also validating the inferred solution.
@@ 5 Discussion
We analyze the impact of the hyperparameters — variance schedule ($\beta_{t}$) and diffusion steps ($T$), on ConDiSim’s performance in Section 5.1, with further discussion on overall scalability, robustness, and reliability in Sections 5.2, 5.4, and 5.5 respectively. This study provides practical insights for optimizing accuracy, efficiency, and reliability across various simulation-based inference tasks.
5.1 Variance Schedule & Diffusion Steps
We conducted an ablation study on the Two Moons test problem to examine the effects of diffusion steps ($T$) and variance schedules ($\beta_{t}$) on posterior approximation accuracy. Varying $T$ from 100 to 1000, we calculate the mean Classifier Two-Sample Test (C2ST) accuracy across five runs with a fixed simulation budget of 10,000 samples, while holding architecture and other hyperparameters constant. The study compared linear, cosine, and quadratic schedules to assess their impact under varying diffusion granularity, with $T$ controlling the refinement steps and $\beta_{t}$ setting the noise rate in the forward diffusion process.
Figure 6: Mean C2ST score across diffusion steps for linear, cosine, and quadratic schedules on the Two Moons benchmark.
@@ 5.1 Variance Schedule & Diffusion Steps
Table 1: Comparison of C2ST scores across SBI benchmarks: (mean $\pm$ std.) at varying simulation budgets.
@@
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.9995 $\pm$0.0007 0.9978 $\pm$0.0016 0.9994 $\pm$0.0004 0.9998 $\pm$0.0001 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 1.0 $\pm$0.0000 NPE 0.9714 $\pm$0.0144 0.9468 $\pm$0.0157 0.9359 $\pm$0.0127 0.9354 $\pm$0.0093 0.9114 $\pm$0.0082 0.9065 $\pm$0.0105 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0000 APT 0.9823 $\pm$0.0055 0.9451 $\pm$0.0158 0.9130 $\pm$0.0301 0.8455 $\pm$0.0116 0.8596 $\pm$0.0152 0.8461 $\pm$0.0135 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0001 Simformer 0.9198 $\pm$0.0451 0.8814 $\pm$0.0433 0.8717 $\pm$0.0453 0.9814 $\pm$0.0106 0.8600 $\pm$0.0452 0.7750 $\pm$0.0473 0.$9$$\pm$0.0003 0.$9$$\pm$0.0012 1.$0$$\pm$0.0002 FMPE 0.9713 $\pm$ 0.0117 0.9779 $\pm$ 0.0095 0.9387 $\pm$ 0.0377 0.9742 $\pm$ 0.0113 0.6661 $\pm$ 0.0307 0.6653 $\pm$ 0.0377 0.$9$$\pm$ 0.0178 0.$9$$\pm$ 0.0345 0.$9$$\pm$ 0.0018 ConDiSim 0.9722 $\pm$0.0368 0.9366 $\pm$0.02395 0.9347 $\pm$0.0368 0.6430 $\pm$0.0237 0.6155 $\pm$0.0317 0.6166 $\pm$0.0480 0.9996 $\pm$0.0002 0.9995 $\pm$0.0002 0.9996 $\pm$0.0002
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7474 $\pm$0.0364 0.6977 $\pm$0.0260 0.7234 $\pm$0.0599 0.9930 $\pm$0.0038 0.9916 $\pm$0.0046 0.9929 $\pm$0.0059 0.9989 $\pm$0.0004 0.9976 $\pm$0.0016 0.9988 $\pm$0.0010 0.9687 $\pm$0.0286 0.9436 $\pm$0.0203 0.9723 $\pm$0.0140 NPE 0.6118 $\pm$0.0155 0.5830 $\pm$0.0240 0.5818 $\pm$0.0158 0.5472 $\pm$0.0076 0.5238 $\pm$0.0112 0.5266 $\pm$0.0037 0.5562 $\pm$0.0147 0.5292 $\pm$0.0068 0.5256 $\pm$0.0066 0.8996 $\pm$0.0206 0.8502 $\pm$0.0174 0.8546 $\pm$0.0213 APT 0.5457 $\pm$0.0058 0.5425 $\pm$0.0124 0.5467 $\pm$0.0050 0.5367 $\pm$0.0038 0.5293 $\pm$0.0052 0.5258 $\pm$0.0047 0.5401 $\pm$0.0115 0.5302 $\pm$0.0069 0.5264 $\pm$0.0075 0.8164 $\pm$0.0295 0.6882 $\pm$0.0342 0.6473 $\pm$0.0343 Simformer 0.5166 $\pm$0.0210 0.5100 $\pm$0.0090 0.5126 $\pm$0.0070 0.6130 $\pm$0.0281 0.6226 $\pm$0.0398 0.5866 $\pm$0.0430 0.6501 $\pm$0.0519 0.6207 $\pm$0.0519 0.5780 $\pm$0.0211 0.7661 $\pm$0.0685 0.6201 $\pm$0.0490 0.7192 $\pm$0.0721 FMPE 0.7894 $\pm$ 0.0254 0.7872 $\pm$ 0.0049 0.5821 $\pm$ 0.0292 0.9605$\pm$ 0.0019 0.9716 $\pm$ 0.0023 0.9708 $\pm$ 0.0025 0.9218 $\pm$ 0.0122 0.9100 $\pm$ 0.0174 0.9082 $\pm$ 0.0178 0.9698 $\pm$ 0.0132 0.9686 $\pm$ 0.0139 0.9686 $\pm$ 0.0139 ConDiSim 0.6602 $\pm$0.0098 0.6181 $\pm$0.0237 0.6233 $\pm$0.0276 0.5809 $\pm$0.0210 0.5656 $\pm$0.0197 0.5505 $\pm$0.0056 0.6572 $\pm$0.0215 0.6468 $\pm$0.0364 0.6448 $\pm$0.0216 0.8691 $\pm$0.0132 0.8324 $\pm$0.0281 0.8210 $\pm$0.0188
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7770 $\pm$0.0630 0.6634 $\pm$0.0503 0.7467 $\pm$0.1036 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.9997 $\pm$0.0005 0.9990 $\pm$0.0015 0.9912 $\pm$0.0161 0.9838 $\pm$0.0157 NPE 0.5749 $\pm$0.0331 0.5931 $\pm$0.0329 0.5411 $\pm$0.0195 0.9121 $\pm$0.0338 0.8988 $\pm$0.0306 0.8898 $\pm$0.0565 0.8378 $\pm$0.0396 0.9122 $\pm$0.0207 0.9405 $\pm$0.0125 APT 0.5189 $\pm$0.0080 0.5202 $\pm$0.0114 0.5147 $\pm$0.0125 0.8404 $\pm$0.0107 0.8328 $\pm$0.0156 0.8353 $\pm$0.0164 0.9433 $\pm$0.0076 0.9434 $\pm$0.0061 0.9457 $\pm$0.0068 Simformer 0.5273 $\pm$0.0229 0.5190 $\pm$0.0174 0.5098 $\pm$0.0084 0.6584 $\pm$0.0381 0.6131 $\pm$0.0225 0.6154 $\pm$0.0265 0.9378 $\pm$0.0053 0.9520 $\pm$0.0064 0.9478 $\pm$0.0049 FMPE 0.8051 $\pm$ 0.0506 0.7745 $\pm$ 0.0261 0.5706 $\pm$ 0.0209 0.9266 $\pm$ 0.0288 0.9118 $\pm$ 0.0376 0.8431 $\pm$ 0.0393 0.9573 $\pm$ 0.0013 0.8454 $\pm$ 0.0014 0.8101 $\pm$ 0.0020 ConDiSim 0.5291 $\pm$0.0143 0.5395 $\pm$0.0248 0.5242 $\pm$0.0063 0.6589 $\pm$0.0421 0.5892 $\pm$0.0100 0.6146 $\pm$0.0371 0.8324 $\pm$0.0332 0.8083 $\pm$0.0157 0.8043 $\pm$0.0138
@@
Table 2: Training time (in minutes) for different methods across simulation budgets (mean $\pm$ std. dev.).
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 117.67 $\pm$2.43 164.90 $\pm$57.10 249.50 $\pm$105.05 118.62 $\pm$0.18 179.69 $\pm$49.96 253.48 $\pm$113.38 227.16 $\pm$1.44 241.24 $\pm$7.58 250.47 $\pm$2.22 NPE 7.04 $\pm$0.21 14.21 $\pm$0.04 20.95 $\pm$0.37 12.17 $\pm$0.17 24.96 $\pm$0.64 37.43 $\pm$0.22 20.10 $\pm$5.45 29.62 $\pm$3.74 44.92 $\pm$9.24 APT 90.31 $\pm$8.99 186.17 $\pm$13.79 268.28 $\pm$32.88 101.39 $\pm$7.89 197.30 $\pm$9.32 267.38 $\pm$21.93 40.15 $\pm$3.90 73.74 $\pm$6.25 112.51 $\pm$2.51 Simformer 3.96 $\pm$0.04 7.36 $\pm$0.11 10.73 $\pm$0.25 17.87 $\pm$0.18 35.05 $\pm$0.52 52.81 $\pm$0.91 4.24 $\pm$0.29 7.67 $\pm$0.38 10.68 $\pm$0.45 FMPE 2.63 $\pm$ 0.05 4.73$\pm$ 1.11 10.655 $\pm$ 0.04 2.65 $\pm$ 0.05 5.75 $\pm$ 1.43 7.64 $\pm$ 1.38 2.78 $\pm$ 0.15 6.39 $\pm$ 0.19 8.94 $\pm$ 2.94 CondiSim 0.84 $\pm$0.28 1.682 $\pm$0.09 2.835 $\pm$0.70 0.31 $\pm$0.11 0.60 $\pm$0.14 0.88 $\pm$0.17 0.87 $\pm$0.16 0.94 $\pm$0.13 1.59 $\pm$0.24
@@
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 74.37 $\pm$2.59 100.59 $\pm$2.39 102.30 $\pm$0.51 114.76 $\pm$1.19 115.22 $\pm$1.43 144.10 $\pm$24.15 78.91 $\pm$6.32 75.73 $\pm$0.70 74.99 $\pm$0.20 NPE 21.92 $\pm$2.41 43.92 $\pm$2.58 64.38 $\pm$7.13 25.39 $\pm$2.09 61.43 $\pm$2.61 85.94 $\pm$6.55 80.67 $\pm$14.61 160.92 $\pm$42.82 188.04 $\pm$29.46 APT 26.30 $\pm$0.36 49.96 $\pm$2.73 74.31 $\pm$2.54 21.87 $\pm$1.28 40.00 $\pm$1.37 60.94 $\pm$3.00 45.52 $\pm$4.85 96.15 $\pm$13.77 144.33 $\pm$18.33 Simformer 15.98 $\pm$1.22 31.98 $\pm$2.34 47.85 $\pm$1.86 16.96 $\pm$0.51 32.46 $\pm$1.89 48.01 $\pm$1.44 5.0 $\pm$0.34 9.23 $\pm$0.54 13.64 $\pm$0.21 FMPE 2.76 $\pm$ 0.08 3.08 $\pm$ 1.76 12.97 $\pm$ 0.87 2.44 $\pm$ 0.44 4.42 $\pm$ 1.37 8.05 $\pm$ 1.38 3.63 $\pm$ 0.42 6.97 $\pm$ 0.03 9.55 $\pm$ 0.72 CondiSim 1.31 $\pm$0.07 1.83 $\pm$0.32 2.08 $\pm$0.34 0.43 $\pm$0.04 0.62 $\pm$0.08 0.93 $\pm$0.17 0.60 $\pm$0.11 1.15 $\pm$0.04 1.41 $\pm$0.08
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 114.07 $\pm$2.11 111.30 $\pm$0.28 137.46 $\pm$19.16 48.19 $\pm$0.24 48.70 $\pm$0.84 49.12 $\pm$0.49 43.27 $\pm$0.36 66.43 $\pm$5.70 69.10 $\pm$0.15 162.72 $\pm$14.73 196.19 $\pm$119.42 99.39 $\pm$5.63 NPE 19.56 $\pm$2.47 42.88 $\pm$4.54 64.03 $\pm$6.51 12.03 $\pm$0.80 24.69 $\pm$1.63 38.79 $\pm$1.23 16.55 $\pm$0.85 30.90 $\pm$2.77 49.42 $\pm$1.49 42.97 $\pm$6.35 97.70 $\pm$6.66 117.61 $\pm$4.12 APT 76.27 $\pm$7.02 167.88 $\pm$15.99 240.13 $\pm$22.74 70.47 $\pm$0.88 135.38 $\pm$5.75 203.27 $\pm$5.97 78.61 $\pm$2.11 158.57 $\pm$9.32 219.44 $\pm$11.83 93.40 $\pm$1.26 150.29 $\pm$11.10 216.30 $\pm$21.38 Simformer 3.84 $\pm$0.04 7.15 $\pm$0.09 10.48 $\pm$0.21 18.18 $\pm$0.13 35.46 $\pm$0.17 55.46 $\pm$1.13 17.89 $\pm$0.23 33.47 $\pm$0.40 51.24 $\pm$0.89 4.23 $\pm$0.01 7.67 $\pm$0.13 11.32 $\pm$0.40 FMPE 2.16$\pm$ 0.08 5.36 $\pm$ 0.78 8.31 $\pm$ 1.32 1.78 $\pm$ 0.05 4.40 $\pm$ 0.78 5.94 $\pm$ 4.31 2.15 $\pm$ 0.22 4.30 $\pm$ 3.86 8.71 $\pm$ 2.12 1.36 $\pm$ 0.14 3.71 $\pm$ 2.57 6.34 $\pm$ 2.33 ConDiSim 0.74 $\pm$0.18 1.77 $\pm$0.41 2.31 $\pm$0.59 0.76 $\pm$0.09 1.95 $\pm$0.59 3.46 $\pm$0.70 0.91 $\pm$0.20 2.34 $\pm$0.37 3.02 $\pm$0.69 1.13 $\pm$0.08 2.61 $\pm$0.80 2.82 $\pm$0.36
As shown in Figure 6, lower $T$ values generally yield better C2ST accuracy, potentially due to reduced cumulative noise, which preserves crucial posterior structure. The quadratic schedule outperforms others, possibly because its gradual noise addition captures complex structures more effectively without early-stage noise overload. Increasing $T$ leads to a plateau and slight decline in accuracy, suggesting diminishing returns and rising instability, with all schedules converging in performance at high $T$ values.

### v2

@@ 4 Experiments
We evaluate ConDiSim on 10 benchmark problems from the sbibm suite [Lueckmann et al., 2021], and two real‑world problems. For each benchmark, we constrain the training to simulation budgets of 10,00010{,}000, 20,00020{,}000, and 30,00030{,}000 simulations, and infer the posterior by drawing 10,00010{,}000 samples from the learned model. We assess posterior quality using Maximum Mean Discrepancy (MMD), the Classifier Two-Sample Test (C2ST) [Friedman, 2003], and Empirical Cumulative Distribution Function (ECDF) diagnostics from the BayesFlow framework [Radev et al., 2023b]. C2ST results are reported in the main text; MMD scores and full ECDF plots for all benchmarks are provided in Appendix F. Complete hyperparameter configurations for all methods are detailed in Appendix C.
@@
(a) SLCP.
(b) SLCP Distractors.
Figure 3: SLCP: Comparison of generated vs. reference posteriors for SLCP and SLCP Distractors.
@@
Figure 4: Posterior Distributions for Hodgkin–Huxley model.
@@
(b) Posterior Predictive Trajectories.
Figure 5: Posterior Distributions for Genetic Oscillator model.
4.1 Benchmark Posterior Analysis
In this section, we analyze three representative benchmarks: Two Moons, SLCP, and SLCP Distractors, selected for their distinct and complementary challenges in posterior inference. The Two Moons benchmark poses the challenge of capturing a bi-modal posterior with a distinctive crescent geometry. As depicted in Figure 2(a), ConDiSim accurately recovers both modes and the complex posterior structure. The ECDF plots for $\theta_{1}$ and $\theta_{2}$ (Figure 2(b)) demonstrate that the inferred distributions are well-calibrated, with the majority of rank statistics falling within the 90% confidence bands; a minor deviation for $\theta_{1}$ suggests a slight degree of overconfidence.
The SLCP benchmark encompasses a range of parameter behaviors including uniform, Gaussian, multi-modal, and skewed distributions. Its variant, SLCP Distractors, further tests the model’s robustness by incorporating uninformative noise dimensions, thereby simulating scenarios with heavily corrupted data. Figures 3(a) and 3(b) illustrate that ConDiSim is able to capture these complex posterior landscapes with high fidelity, and also effectively isolate the informative parameters from the distractors. The rank ECDF plots for both SLCP variants (Figures 14(a) and 14(b) in the Appendix) also remain centered within the confidence intervals, confirming that the model reliably captures parameter variability even under noisy conditions. Detailed descriptions of these benchmarks, are provided in Appendix B.
Table 1 reports C2ST scores across the 10 benchmarks. On tasks such as Two Moons, ConDiSim achieves performance comparable to Simformer, APT, and NPE, while requiring substantially shorter training times (Table 2). Although ConDiSim does not always attain the top C2ST score, the method offers a favorable trade-off between efficiency and posterior quality, which is crucial under limited simulation budgets. Notably, the SLCP task remains challenging for all methods, highlighting the inherent difficulty of accurately recovering high-dimensional, multi-modal posteriors. Moreover, C2ST can be overly sensitive to small errors in individual dimensions, making it less reliable in high-dimensional settings. For example, as shown in Figure 3(b), ConDiSim closely matches the true posterior in most dimensions, with deviations concentrated in the final parameter. To provide additional insight, we report the MMD metric for each benchmark problem in the Supplementary Table 4.
@@ 4.2 Application to Real-World Scenarios
We evaluate the performance of ConDiSim on two real-world simulation models. First, the Hodgkin-Huxley (HH) model [Gloeckler et al., 2024] which simulates neuronal action potentials by modeling the membrane voltage dynamics governed by sodium, potassium, and leak currents. Second, we consider a Genetic Oscillator model which is a relatively high-dimensional gene regulatory network that exhibits oscillatory behavior, driven by a positive-negative feedback loop mimicking a circadian clock [Vilar et al., 2002]. The model encompasses 9 species undergoing 18 reactions, parameterized by 15 reaction rate constants. Although the full system involves multiple interacting components, we focus on the inference of only three species (C, A, R). Detailed description is available in Appendix B.
Unlike the SBI benchmarking problems, the true posterior distributions for these real-world models are unknown. To assess ConDiSim, we use well motivated parameter combinations from literature to generate corresponding synthetic observations (e.g., parameter combination where the genetic oscillator results in stable oscillations). Performing inference based on these observations enables us to evaluate the inferred posterior by comparing its concentration around the true parameter value. We also perform a posterior predictive check to validate the model. In this procedure, five posterior samples are randomly selected and used to simulate corresponding observations, which are then compared with the observed data. This check provides an additional measure of the consistency and reliability of the inferred posteriors.
For the HH model, Figure 4(a) shows that ConDiSim effectively infers a posterior distribution concentrated around the true parameter values, while Figure 4(b) confirms that the simulated voltage trajectories closely match the reference signals. The parameter estimation task for the genetic oscillator is more challenging due to broad and varied parameter priors. Figure 5(a) indicates that the peak of the posterior distribution is near the true parameter values. Figure 5(b) illustrates both the true and generated trajectories for three selected species, obtained by randomly sampling three posterior sets and plotting the corresponding simulations. The trajectories align well with the true data, with slight deviations emerging over time due to the inherent stochastic nature of the model, as expected. Figure 6 (Appendix) depicts multiple trajectories simulated using the true parameter combination - which closely match the inferred results. In both cases, the trajectories are very close to each other during initial time steps, and develop small deviations owing to model stochasticity as time steps evolve. The plots highlight the system’s inherent variability, and also validate the inferred solution. From Figs.5(a), 9, we also see that ConDiSim is able to infer a more accurate, tighter posterior as compared to Simformer, highlighting its suitability for challenging inference problems involving stochastic, discrete-time observations.
@@ 5 Discussion
In this section, we compare ConDiSim against established SBI methods (Section 5.1) and analyze its behavior along three practical axes: scalability with problem dimensionality (Section 5.2), robustness to simulator noise or misspecification (Section 5.3), and reliability in capturing posterior uncertainty (Section 5.4). This analysis provides practical guidance for deploying diffusion-based inference methods by identifying regimes where ConDiSim excels and scenarios where alternative approaches may be more suitable.
5.1 ConDiSim vs. Other SBI Methods
Tables 1 and 4 show that diffusion-based methods (Simformer, ConDiSim) often achieve higher posterior accuracy than flow-based (NPE, APT, FMPE) and GAN-based (GATSBI) approaches, particularly on problems with multimodal or highly nonlinear posteriors. Simformer attains the best overall performance on several benchmarks (Table 4), owing to its transformer architecture with cross-attention conditioning; however, this comes at the cost of substantially longer training times (Table 2). In contrast, ConDiSim uses a lightweight FiLM-based conditioning mechanism that avoids costly attention computations, achieving competitive accuracy with significantly lower training overhead - while its iterative denoising incurs no practical slowdown in inference. NPE and APT remain strong baselines on simpler or lower-dimensional tasks, where their one-shot or sequential flow-based strategies are sufficient. GATSBI, meanwhile, is hampered by the well-known instabilities of adversarial training. We also note from Table 1 that ConDiSim outperforms other methods on problems involving noisy, discrete observations, such as Bernoulli GLM, Bernoulli GLM Raw and SIR, leveraging FiLM. This makes the proposed approach well-suited for SBI problems in computational biology where parameter inference of stochastic models with time-series observations is often encountered (e.g., the genetic oscillator test problem, as noted before).
@@
Table 1: C2ST scores across SBI benchmarks (mean $\pm$ std.) at varying simulation budgets; best in bold, second-best underlined.
@@
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 0.9714 $\pm$0.0144 0.9468 $\pm$0.0157 0.9359 $\pm$0.0127 0.9354 $\pm$0.0093 0.9114 $\pm$0.0082 0.9065 $\pm$0.0105 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0000 APT 0.9823 $\pm$0.0055 0.9451 $\pm$0.0158 0.9130 $\pm$0.0301 0.8455 $\pm$0.0116 0.8596 $\pm$0.0152 0.8461 $\pm$0.0135 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0001 GATSBI 0.9995 $\pm$0.0007 0.9978 $\pm$0.0016 0.9994 $\pm$0.0004 0.9998 $\pm$0.0001 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 1.0 $\pm$0.0000 Simformer 0.9198 $\pm$0.0451 0.8814 $\pm$0.0433 0.8717 $\pm$0.0453 0.9814 $\pm$0.0106 0.8600 $\pm$0.0452 0.7750 $\pm$0.0473 0.9999 $\pm$0.0003 0.9999 $\pm$0.0012 1.0000 $\pm$0.0002 FMPE 0.9713 $\pm$ 0.0117 0.9779 $\pm$ 0.0095 0.9387 $\pm$ 0.0377 0.9742 $\pm$ 0.0113 0.6661 $\pm$ 0.0307 0.6653 $\pm$ 0.0377 0.9999 $\pm$ 0.0178 0.9999 $\pm$ 0.0345 0.9999 $\pm$ 0.0018 ConDiSim 0.9456 $\pm$0.0245 0.9222 $\pm$0.02609 0.8832 $\pm$0.08519 0.6673 $\pm$0.0668 0.6099 $\pm$0.0272 0.6041 $\pm$0.0237 0.9995 $\pm$0.0001 0.9995 $\pm$0.0005 0.9995 $\pm$0.0003
[…]
@@
Table 2: Training time (in minutes) for different methods across simulation budgets (mean $\pm$ std. dev.); best in bold, second-best underlined.
[…]
@@
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 21.92 $\pm$2.41 43.92 $\pm$2.58 64.38 $\pm$7.13 25.39 $\pm$2.09 61.43 $\pm$2.61 85.94 $\pm$6.55 80.67 $\pm$14.61 160.92 $\pm$42.82 188.04 $\pm$29.46 APT 26.30 $\pm$0.36 49.96 $\pm$2.73 74.31 $\pm$2.54 21.87 $\pm$1.28 40.00 $\pm$1.37 60.94 $\pm$3.00 45.52 $\pm$4.85 96.15 $\pm$13.77 144.33 $\pm$18.33 GATSBI 74.37 $\pm$2.59 100.59 $\pm$2.39 102.30 $\pm$0.51 114.76 $\pm$1.19 115.22 $\pm$1.43 144.10 $\pm$24.15 78.91 $\pm$6.32 75.73 $\pm$0.70 74.99 $\pm$0.20 Simformer 15.98 $\pm$1.22 31.98 $\pm$2.34 47.85 $\pm$1.86 16.96 $\pm$0.51 32.46 $\pm$1.89 48.01 $\pm$1.44 5.0 $\pm$0.34 9.23 $\pm$0.54 13.64 $\pm$0.21 FMPE 2.76 $\pm$ 0.08 3.08 $\pm$ 1.76 12.97 $\pm$ 0.87 2.44 $\pm$ 0.44 4.42 $\pm$ 1.37 8.05 $\pm$ 1.38 3.63 $\pm$ 0.42 6.97 $\pm$ 0.03 9.55 $\pm$ 0.72 CondiSim 2.02 $\pm$0.27 5.16 $\pm$0.84 8.44 $\pm$1.02 3.22 $\pm$0.72 7.22 $\pm$0.50 10.67 $\pm$0.69 2.42 $\pm$0.21 7.19 $\pm$0.97 11.19 $\pm$2.31
[…]